A company manufactures 2,000 units; calculate the average production cost with $35 per unit. In a year, production goes down to 1,800 units. What happens to the average cost per unit?
Answer #1

Explaination Aoubt Average Production cost-

Average cost is the total amount of all production costs divided by the quantity of output produced. This number is also known as average total cost or unit cost. In simpler terms, it measures how much a business has to spend on each unit or product of output produced. You can determine the ATC with a simple equation:

Average Total Cost = Total Cost of Production / Quantity of Units Produced

How change-

Average cost includes fixed costs, like those necessary for production, that remain the same no matter the output. An example of a fixed cost is the building space and equipment used to assemble a product. Average cost also includes variable costs. Examples of variable costs are specific parts needed to build a product, which may increase or decrease according to output.

Understand with example:-

A company produce 1000 units and its fixed cost is $1000 and variable cost is $1/unit then APC will be= $(1000+1000*1)/1000 units= $2/unit

now the company reduce its production to 500 units then
APC= $(1000+1*500)/500units=$3/unit

now the company reduce its production to 2000 units then
APC=$(1000+1*2000)/2000 units=1.5unit

conclusion:- APC (Average production cost) include both componant Fixed as well as variable, with the increase or decrease in units,variable cost varry accourding. but the fixed componant will remain the same. if the Number of units is increase then the allocation of the fixed cost over the units will less so average production cost decrease.(1000/2000=$.5/unit)


On the other side if the Number of units production decrese then allocation of the fixed cost will be more and over the less units(1000/500=$2/unit) so average production cost will increase.

Result;- In the given case production unit is decrease to 1800 the accordingly the APC will increase as per the concept above discussed.